OpenStudy (anonymous):

Need help find three points that solve the eqaution the plot them in the graph -3x+2y=2 thanks

OpenStudy (amistre64):

let x=0, let y=0, let x=1 ...

OpenStudy (anonymous):

Huh

OpenStudy (amistre64):

youll have to be more clear about your confusion

OpenStudy (anonymous):

-3x+2y=2 \[2y=2+3x\] \[y=\frac{ 2+3x }{2 }\] when x=0, substitute x=0+2=2 \[y=\frac{ 2+3\left( 2 \right) }{2 }=\frac{ 8 }{2 }=4\] every time you add successively 2 to x ,we get a new value of y e.g.,x=4,6,8............ y=(2+3x)/2 got it.